### v2

@@ Abstract
Over the past six years, molecular transformer models have become part of the computational toolbox for drug design tasks. Most of the current transformer-based chemical language models are pre-trained on millions to billions of molecules from unlabeled datasets like ZINC, or ChEMBL. However, the improvement from such scaling in dataset size is not confidently linked to improved molecular property prediction.
The aim of this study was to investigate and overcome some of the limitations of transformer models in predicting molecular properties. Specifically, we examine the impact of pre-training dataset size and diversity on the performance of transformer models, and investigate the use of domain adaptation as a technique for improving model performance. First, our findings indicate that increasing pre-training dataset size beyond $\sim 400K$ molecules (i.e., 30%) from the pre-training dataset does not result in a significant improvement on four ADME endpoints, namely, solubility, permeability, microsomal stability, and plasma protein binding. Further analysis indicates that the limited benefits of increasing the pre-training dataset size may be attributed to its lower diversity in Murcko scaffolds compared to the downstream datasets. Second, our results demonstrate that using domain adaptation by further training the transformer model on a small set of domain-relevant molecules, i.e., a few hundred to a few thousand, using multi-task regression of physicochemical properties was sufficient to significantly improve performance for three out of the four investigated ADME endpoints (P-value $<0.001$). Finally, we observe that a model pre-trained on $\sim 400$K molecules and domain adopted on a few hundred/thousand molecules perform similarly (P-value $>0.05$) to more complicated transformer models like MolBERT [1] (pre-trained on $\sim 1.3$M molecules) and MolFormer [2] (pre-trained on $\sim 100$M molecules). A comparison to a random forest model trained on basic physicochemical properties showed similar performance to the examined transformer models.
We believe that current transformer models can be improved through further systematic analysis on pre-training and downstream data, pre-training objectives, and scaling laws, ultimately leading to better and more helpful models.
@@ 1 Introduction
Although current molecular transformer models have been pre-trained on millions to billions of molecules, investigations have shown that increasing the pre-training dataset size does not consistently lead to improved predictions of molecular properties [15]. Several studies have explored the impact of increasing pre-training dataset size [2, 22, 27, 28], but this has not yet been done exhaustively, and these studies often lack distribution or significance analysis, which is crucial for ensuring the robustness. For example, pre-training on molecules combined from different databases, like ZINC and PubChem as in MolFormer [2] or ZINC, ChEMBL, and PubChem as in Chen et al. [28], did not show noticeable differences in performance compared to using a single database. Additionally, it has recently been demonstrated in the field of material science that large databases often contain a substantial amount of redundant information [29], while for the protein language modeling field, a small but diverse pre-training dataset was sufficient to improve performance on downstream analysis [30].
@@
Our findings from the six datasets of the ADME benchmark [5] indicate that increasing pre-training dataset size was either not beneficial or only helpful up to a certain fraction of the dataset (i.e., $\ll 1$M molecules). We also observed that applying domain adaptation using the MTR objective significantly improved the model’s performance for solubility, permeability, and microsomal stability endpoints by showing P-values < 0.01, as well as noticeably increasing the average $R^{2}$ by more than 0.1 unit (i.e., large effect size). Additionally, transformer models can be trained much more efficiently than existing models while preserving or even enhancing predictive power. Finally, all examined transformer models performed similarly to a basic RF model with default settings on a pre-defined physicochemical properties as a molecular representation on the examined datasets.
We believe that molecular transformer models are yet to be optimized to address the specific needs and nuances of molecular data, such as ensuring diversity in the pre-training dataset, selecting the most suitable training objectives, and understanding the relation between upstream and downstream datasets. Therefore, we expect these models to surpass current baselines when such measures are taken into account.
@@ 2 Datasets and Preprocessing
In this work, we pre-train a transformer model on a dataset sampled from a large-scale general-purpose library of molecules, namely GuacaMol [18]. We then evaluate our models on a benchmark dataset that explores ADME properties (Absorption, distribution, metabolism, excretion) [5]. In our approach, datasets are utilized in three stages of model training: pre-training, domain adaptation, and evaluation. The GuacaMol dataset is exclusively used for pre-training, the unlabeled ADME dataset for domain adaptation, and the labeled ADME dataset version for evaluation. In this section, we provide a brief description of each dataset.
@@ 2.1 Pre-training Dataset: GuacaMol
2.1.1 Pre-training dataset selection
Besides training a model on the full training split of the GuacaMol [18] dataset (i.e., $\sim 1.3$M molecules), we investigate three more models trained on 0% (i.e., randomly initialized model with no pre-training), 30%, and 60% of the datasets. To select the 30% and 60% subsets, the data was first clustered using BitBirch [32] based on Morgan fingerprints (radius 2, 2084 bits) and tanimoto similarity, then selections were made to ensure the same cluster distribution per subset. RDKit [33] was used to calculate fingerprints and similarity, and the selection procedure was adopted from talktorial T005 from TeachOpenCADD [34].
@@ 2.2 Domain Adaptation and Downstream Datasets: ADME benchmark
The domain adaptation (DA) step is introduced to further train the pre-trained model on a smaller domain-specific dataset. In our analysis, the molecules of the downstream dataset are passed to the pre-trained model without labels for DA. Later, in the MPP task the ADME dataset molecules including their labels are used for evaluation. The full benchmark dataset contains $\sim 3.5$K molecules, however, each of the four ADME endpoints was tested on a subset only. Table 1 shows the number of molecules tested per endpoint. In the following, we will detail the different endpoints.
# Molecules # Molecules without censored labels
Solubility 2173 2173
Permeability 2642 2642
HLM 3087 2129
RLM 3054 2708
hPPB 194 194
rPPB 168 168
Table 1: Number of molecules per endpoint with and without censored labels. The numbers in bold refer to the datasets that changed in size after removing censored labels.
Solubility is a vital property for understanding the uptake and distribution of a molecule within living organisms, which determines a molecule’s efficacy and usability [35]. Solubility assays inform about maximum amount of a molecule to dissolve in a solution, either water or a physiologically relevant medium, under certain pH and temperature [36]. Fang et. al. [5] measured the solubility of $\sim 2.2$K molecules in phosphate buffered saline (PBS) at pH 6.8 following the protocol in Kestranek et. al. [37]. The reported measuring unit is $\mu g/mL$. Figure1 shows the distribution of the measurements. The data has a bimodal distribution with a major peak at $\sim 1.5$ log($\mu g/mL$) ($\sim 31.62$ $\mu g/mL$) and another small peak at $\sim-0.5$ log($\mu g/mL$) ($\sim 0.316$ $\mu g/mL$).
@@
Permeability assays determine the rate at which a molecule passes through a biological membrane by measuring the ratio between movement to and from the bloodstream [38, 39]. Fang et. al. uses the MDR1 transfected MDCK cell lines to measure this efflux ratio [5]. MDR1 (also known as P-glycoprotein or P-gp receptor) is present in various tissues, such as intestines and blood-brain barrier, making it a good candidate to assess permeability [40, 38, 39]. Figure 1 shows the distribution of the $\sim 2.6$K molecule’s efflux ratio, where $\sim 50\%$ of the data is concentrated at $\sim-0.2$ Log unit (an efflux ratio of $\sim 0.6$).
@@
Microsomal stability is used to measure the clearance of a molecule from the body, specifically, intrinsic clearance $CL_{int}$ [41]. Intrinsic clearance quantifies the volume of incubation medium from which a molecule is cleared per time unit per microsomal weight unit. In Fang et. al. [5], the measuring unit is mL/min/kg presented for two datasets: human liver microsomes (HLM) and rat liver microsomes (RLM). Figure 1 shows the data summary of the two datasets with HLM spanning around three log units of mL/min/kg (0.6 - 3.4), while RLM spans around four log units (1 - 4). The HLM dataset has a sharp peak at $\sim 0.67$ ($\sim 4.7$ mL/min/kg). Interestingly, the boxplot and the scatter plot of the individual data points show that $\sim 27\%$ of the data has the same target value. This observation corresponds to what is called "censored labels", where experimental values can be recorded only up to a certain threshold, and every observation beyond this threshold is set to a constant value [42]. When training a regression model to predict continuous values, this censoring is introducing noise to the data. The same censoring problem is seen for the RLM dataset, although less severely as only $\sim 9$% of the data is censored. While the censored regression problem has been investigated in fields like survival analysis, its resolution in MPP is still in progress [42]. In this work, we handle the censored label by removing the molecules with censored labels from the evaluation step to reduce the noise. However, we keep the molecules during the DA step as they still provide information for further training. After removing these values, the number of labeled molecules becomes $\sim 2.1$K and $\sim 2.7$K for HLM and RLM datasets, respectively. Suppl. Figure S1 and Table 1 show the new distribution and summary of the two datasets. The HLM dataset now spans the range 0.7 to 3 log units while RLM spans the same range (1 to 4 log units).
Plasma protein binding assays measure the amount of a compound that has not been captured by the plasma proteins in the blood. This assays is important because it is believed that the percentage of free molecules in the blood is responsible for a compound’s activity and efficacy. Therefore, knowing the percentage of free molecules can also guide other assays like permeability and clarence [43]. Fang et. al reported the percentage of unbound molecules in the blood in both human (hPPB) and rats (rPPB). Both datasets contained very few molecules ($<200$ molecules) and the measurements span the range -2 to 2 Log unit (0.01 - 100% unbound molecules), see Figure1.
Figure 1: Data summary and distributions of the six investigated endpoint.
2.3 Dataset Splitting
For pre-training, we used the specified split of the GuacaMol dataset into train and test splits with an 80:20 ratio. For domain adaptation, all endpoint molecules were used for further-training. Finally, 5-fold cross validation (CV) was performed for evaluating the predictions on the labeled molecules. We report two types of splitting for CV: random splitting, and cluster-based splitting using DataSAIL [44] with default settings. To avoid data leakage, DataSAIL splits data by minimizing the similarity between any two splits formulated as a Binary Linear Program (BLP) to optimize the solution. 5-fold cluster-based splitting was possible for all datasets except for hPPB, which could only be split 3-fold.
@@ 3.2 Training objectives
The transformer models learn by optimizing a self-supervised training objective like masked language modeling (MLM). In this work, we used MLM as an objective for pre-training, and explored it for domain adaptation as well. Multi-task regression (MTR) was used for domain adaptation and as a pre-training objective for one of the experiments. Contrastive learning (CL) was used for domain adaptation only. below, we briefly explain each of these objectives. For all objectives, we used a batch size of 16 with a learning rate of $3e^{-5}$ for 20 epochs. The maximum input length was 128 and mean pooling of the tokens was applied to extract a molecule’s representation.
@@ 3.3 Models benchmarking
To identify the strengths and weaknesses of our model, we compare the performance in downstream task prediction when training a RF model using embeddings of our models versus embeddings of two other transformer models, MolBERT [1] and MolFormer [2], as well as the physicochemical properties calculated by RDKit [33]. We used the provided MolBERT [1] implementation and the MolFormer’s huggingface [50] implementation to extract the embeddings. The same RF model with default settings as provided by scikit-learn [48] was used to predict the molecular properties using the embeddings of our models, MolBERT, MolFormer, and the physicochemical properties. Table 2 shows the differences between all models examined. The RF + PhysChem model is considered as a simple and transparent baseline model to compare against the remaining models.
@@
Table 2: An overview of the benchmarked models explaining their training decision in terms of pre-training dataset size, objectives and number of trainable parameters. MLM: masked language modeling, MTR: multi-task regression of physicochemical properties, SMILES-Eq: predicting whether two SMILES strings correspond to the same molecule, and CL: contrastive learning of the different representations of the SMILES string.
3.4 Evaluation metrices
During training, we normalize the response values using the standard scaling, for the evaluation we transform predictions back to the original values. We report performances using the coefficient of determination ($R^{2}$, equation 6), which measures the goodness of fit by comparing the model’s residual sum of squares (RSS, equation 4) (i.e., true value $y_{i}$ to predicted value $\hat{y_{i}}$) to the total sum of squares (TSS, equation 5) (i.e., true value $y_{i}$ to the mean value $\bar{y}$). A value of zero means that the model is performing as good as a model that only predicts the mean value of a dataset. A value $>0$ corresponds to a model that gives better prediction than such mean-only model, while a value $<0$ corresponds to a model worse than the mean-only model. The maximum value of the metric is 1 (i.e., the model predicted each value correctly) and the lowest value is $\infty$- ∞. The sklearn package [48] was used to calculate the $R^{2}$ metric.
$\text{RSS}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (4)
$\text{TSS}=\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ (5)
$R^{2}=1-\frac{\text{RSS}}{\text{TSS}}$ (6)
In our analysis, we show the $R^{2}$ of the five folds as either a boxplot or an error barplot using the standard error of means (SEM, equation 7), which assesses the variability of model estimates using the sample standard error $\sigma$ and the sample size $n$.
$\text{SEM}=\frac{\sigma}{\sqrt{n}}$ (7)
For significance analysis, we report P-values comparing the different models using the non-parametric Mann-Whitney U test (also known as Wilcoxon rank-sum test) [51]. This test does not require the data to be normally distributed, and is calculated by ranking the values in two distributions to determine if there is a significant difference in the sum of ranking of each distribution. The P-value of multiple comparisons is corrected using the Holm’s method [52] to account for the problem of increased probability of false positives [53]. The scikit-posthoc package [54] was used to calculate the corrected P-values, and a P-value $<0.05$ was considered significant.

### v3

@@ Abstract
Over the past six years, molecular transformer models have become an integral part of the computational toolbox for drug discovery. Most existing models are pre-trained on millions to billions of molecules from large-scale unlabeled datasets such as ZINC or ChEMBL. However, the extent to which such large-scale pre-training improves molecular property prediction remains unclear.
This study investigates the potential of transformer models for molecular property prediction while addressing their current limitations. We explore strategies to enhance performance, including the influence of pre-training dataset size and the benefits of domain adaptation through chemically informed objectives. Our results show that increasing the pre-training dataset beyond approximately 400K–800K molecules does not improve performance across seven datasets covering five ADME endpoints: lipophilicity, permeability, solubility (two datasets), microsomal stability (two datasets), and plasma protein binding. In contrast, applying domain adaptation on a small number of domain-relevant molecules ($\leq 4K$) using multi-task regression of physicochemical properties significantly improves model performance across all datasets (P-value < 0.001).
Furthermore, we find that a model pre-trained on $\sim$400K molecules and adapted on a small domain-specific dataset outperforms larger-scale transformer models like MolFormer and performs comparably to MolBERT. Benchmarking these models alongside Random Forest (RF) baselines using physicochemical descriptors and Morgan fingerprints reveals that incorporating chemically and physically informed features consistently leads to superior performance, regardless of whether used with traditional or transformer-based architectures.
While traditional models, such as RF, remain strong baselines, this study identifies concrete practices that significantly enhance the performance of transformer models. In particular, aligning pre-training and adaptation with chemically meaningful tasks and domain-relevant data offers a promising path forward for future advancements in molecular property prediction.
Our models are available on HuggingFace to allow for easy use and adaptation at https://huggingface.co/collections/UdS-LSV/domain-adaptation-molecular-transformers-6821e7189ada6b7d0a5b62d4.
@@ 1 Introduction
Although current molecular transformer models have been pre-trained on millions to billions of molecules, investigations have shown that increasing the pre-training dataset size does not consistently lead to improved predictions of molecular properties [13]. Several studies have explored the impact of increasing pre-training dataset size [26, 21, 27, 28], but this has not yet been done exhaustively, and these studies often lack distribution or significance analysis, which is crucial for ensuring the robustness. For example, pre-training on molecules combined from different databases, like ZINC and PubChem as in MolFormer [26] or ZINC, ChEMBL, and PubChem as in Chen et. al.. [28], did not show noticeable differences in performance compared to using a single database. Additionally, it has recently been demonstrated in the field of material science that large databases often contain a substantial amount of redundant information [29], while for the protein language modeling field, a small but diverse pre-training dataset was sufficient to improve performance on downstream analysis [30].
@@
Our experiments across seven datasets spanning five ADME endpoints [3, 31] reveal that increasing the size of the pre-training dataset provided limited benefits, with improvements plateauing at 400 - 800K molecules (30-60% of the GuacaMol dataset [16]). In addition, applying domain adaptation with the MTR objective led to significant performance gains across all datasets (P-values < 0.01), an improvement that was not possible by data scaling alone. Furthermore, our transformer models achieved competitive or superior predictive accuracy compared to existing large scale transformer models(MolFormer [26] and MolBERT [20]), while being more computationally efficient. Notably, models that incorporate chemically informed features and objectives demonstrated the strongest performance, both among transformers and baselines, highlighting the importance of domain-aware training strategies.
@@ 2 Datasets and Preprocessing
In this work, we pre-train a transformer model on a dataset sampled from a large-scale general-purpose library of molecules, namely GuacaMol [16]. We then evaluate our models on a benchmark dataset that explores ADME properties (Absorption, distribution, metabolism, excretion) [3, 31]. In our approach, datasets are utilized in three stages of model training: pre-training, domain adaptation, and evaluation. The GuacaMol dataset is exclusively used for pre-training, the unlabeled ADME dataset for domain adaptation, and the labeled ADME dataset version for evaluation. In this section, we provide a brief description of each dataset.
@@ 2.2 Domain Adaptation and Downstream Datasets: ADME benchmark
For domain adaptation (DA) and evaluation, we utilize seven datasets spanning five ADME endpoints: lipophilicity, permeability, solubility (two datasets), intrinsic clearance (two datasets), and plasma protein binding. The DA step involves further training of a pre-trained model on smaller, domain-specific datasets (i.e., the investigated endpoints) without labels, allowing the model to adapt to the target distribution. Subsequently, in the molecular property prediction (MPP) task, the same datasets—now with labels—are used for evaluation.
The datasets originate from two sources: Fang et. al.. [3] and a ChEMBL collection compiled by AstraZeneca [31]. Fang et. al.. provide datasets for permeability, solubility, two different assay measurements for intrinsic clearance, and two different assay measurements for plasma protein binding. Figure S1 shows the distribution of target values in these datasets. We excluded the plasma protein binding datasets from Fang et. al.. due to their small sample size (fewer than 200 molecules), which is considered insufficient for drawing robust conclusions in regression modeling [33].
The AstraZeneca collection encompasses a wider range of ADME endpoints, including lipophilicity, solubility, three different assay measurements for intrinsic clearance, five different plasma protein binding assays, and pKa measurements for the first and second acidic groups, and the first through third basic groups (see Figure S2). From this collection, we include datasets for lipophilicity, solubility, and one plasma protein binding assay — specifically those containing more than 1,500 molecules, a threshold selected to align with the sample sizes of the Fang et. al.. datasets. Refer to Figure 1 for an overview of the final selected datasets. In the following, we detail each endpoint used in this study.
Figure 1: Data summary and distributions of the seven investigated ADME endpoints. The boxplots are shown per histogram to highlight the location of the different quartiles and datapoints that would be seen as outliers. The plots with two distributions show the presence of censored labels which are not suitable for being used directly in a regression model. Therefore, the two distributions show the original data (blue) and the actual data (orange) used during evaluation.
Lipophilicity is a key physicochemical property describing the affinity of a compound for lipophilic (nonpolar) versus hydrophilic (polar) environments, often expressed as a distribution coefficient (logD) or partition coefficient (logP) [34]. It plays a critical role in drug discovery by influencing absorption, distribution, metabolism, excretion, and toxicity (ADMET) properties [35]. Experimental determination of lipophilicity is commonly performed using the shake flask method, particularly for logD at physiological pH of 7.4. In the dataset sourced from ChEMBL (CHEMBL3301361) [31], lipophilicity data were measured for $\sim$4K molecules by AstraZeneca using the octan-1-ol/water shake flask method, as described by Wenlock et. al.. (2011), covering a logD range of -1.5 to 4.5 [36].
@@
Permeability assays determine the rate at which a molecule passes through a biological membrane by measuring the ratio between movement to and from the bloodstream [37, 38]. Fang et. al. uses the MDR1 transfected MDCK cell lines to measure this efflux ratio [3]. MDR1 (also known as P-glycoprotein or P-gp receptor) is present in various tissues, such as intestines and blood-brain barrier, making it a good candidate to assess permeability [39, 37, 38]. Figure 1 shows the distribution of these $\sim 2.6$K molecule’s efflux ratio.
Solubility is a vital property for understanding the uptake and distribution of a molecule within living organisms, which determines a molecule’s efficacy and usability [40]. Solubility assays inform about maximum amount of a molecule to dissolve in a solution, either water or a physiologically relevant medium, under certain pH and temperature [41]. Fang et. al.. [3] measured the solubility of $\sim 2.2$K molecules in phosphate buffered saline (PBS) at pH 6.8 following the protocol in Kestranek et. al.. [42]. The reported measuring unit is $\mu g/mL$. Figure 1 shows the distribution of the measurements. The data has a strong skew (i.e., imbalance) at $\sim 1.5$ log($\mu g/mL$) ($\sim 31.62$ $\mu g/mL$). The second solubility dataset, measured by AstraZeneca, is obtained from ChEMBL [31]. In this assay, solubility of $\sim 1.8$K molecules was determined in pH 7.4 buffer using solid starting material, following the shake flask method described by Bevan and Lloyd [43]. The experimental concentrations range from 2 to 6 log(nM) (i.e., 0.10 to 1500 $\mu$M). The data is quite uniform for the range 3.5 - 6.2, and underrepresented for the values from 2 - 3.5.
@@
Microsomal stability is used to measure the clearance of a molecule from the body, specifically, intrinsic clearance $CL_{int}$ [44]. Intrinsic clearance quantifies the volume of incubation medium from which a molecule is cleared per time unit per microsomal weight unit. In Fang et. al. [3], the measuring unit is mL/min/kg presented for two datasets: human liver microsomes (HLM) and rat liver microsomes (RLM). Figure 1 shows the data summary of the two datasets with HLM spanning around three log units of mL/min/kg (0.6 - 3.4), while RLM spans around four log units (1 - 4). The HLM dataset has a sharp peak at $\sim 0.67$ ($\sim 4.7$ mL/min/kg). However, further inspection shows that $\sim 27\%$ of the data has the same target value. This observation corresponds to what is called "censored labels", where experimental values can be recorded only up to a certain threshold, and every observation beyond this threshold is set to a constant value [45]. When training a regression model to predict continuous values, this censoring is introducing noise to the model. The same censoring problem is seen for the RLM dataset, although less severely as only $\sim 9$% of the data is censored. While the censored regression problem has been investigated in fields like survival analysis, its resolution in MPP is still in progress [45]. In this work, we handle the censored label by removing the molecules with censored labels from the evaluation step to reduce the noise. However, we keep the molecules during the DA step as they still provide information for further training. After removing these values, the number of labeled molecules becomes $\sim 2.1$K and $\sim 2.7$K for HLM and RLM datasets, respectively. The HLM dataset now spans the range 0.7 to 3 log units while RLM spans the same range (1 to 4 log units).
Plasma protein binding assays measure the amount of a compound that has been captured by the plasma proteins in the blood. This assay is important because the percentage of free molecules in the blood is partly responsible for a compound’s activity and efficacy. Therefore, knowing the percentage of free molecules can also guide other assays like permeability and clearance [46]. A human plasma protein binding (PPB) dataset from AstraZeneca, available in ChEMBL, was measured using equilibrium dialysis. In this assay, compounds are incubated with whole human plasma at 37C for over 5 hours to reach binding equilibrium. The method follows the protocol described by Testa et. al. [47]. For $\sim 1,600$ molecules, experimental binding values range from 10% to 99.95%, equivalent to 0–2 log units. However, the data distribution shows that $\sim$ 70% of the data is concentrated in 0.06 log units (between 70% and 100% bound).
@@ 3.2 Training objectives
The transformer models learn by optimizing a self-supervised training objective like masked language modeling (MLM). In this work, we used MLM as an objective for pre-training, and explored it for domain adaptation as well. Multi-task regression (MTR) was used for domain adaptation and as a pre-training objective for one of the experiments, while contrastive learning (CL) was used for domain adaptation only. Below, we briefly explain each of these objectives. For all objectives, we used a batch size of 16 with a learning rate of $3e^{-5}$ for 20 epochs. The maximum input length was 128 and mean pooling of the tokens was applied to extract a molecule’s representation.
@@
3.3 Pre-training dataset selection
Besides training a model on the full training set of the GuacaMol [16] dataset (i.e., $\sim 1.3$M molecules), we investigate three more models trained on 0% (i.e., randomly initialized model with no pre-training), 30%, and 60% of the GuacaMol dataset. To select the 30% and 60% subsets, the data was first clustered using BitBirch [54] with default settings, and tanimoto similarity based on Morgan fingerprints. Selections were made proportionally from each cluster to preserve the overall cluster distribution across subsets. RDKit [52] was used to calculate fingerprints and similarities, and the selection procedure was adopted from talktorial T005 from TeachOpenCADD [55].
3.4 Dataset splitting and cross validation
For pre-training, we used the predefined split of the GuacaMol dataset, with an 80:20 ratio for training and testing. During domain adaptation, all molecules from the individual ADME datasets were used for further training without labels. For evaluation on labeled data, we followed the recommendations of Ash et. al.. [56], performing $5\times 5$ repeated cross-validation using both random splitting and cluster-based Butina splitting with Morgan fingerprints (as used in Ash et. al.. This cross-validation strategy is designed to produce a large number of independent estimates of the evaluation metric. According to the central limit theorem (CLT), the average of many such estimates tends to follow a normal distribution, enabling more reliable statistical analysis. Satisfying the CLT assumptions allows for the application of parametric statistical tests to assess confidence intervals and significance. In this work, we primarily report results obtained using Butina cluster splitting and refer to results from random splitting where relevant.
3.5 Evaluation metrics
For rigorous evaluation, we again follow the recommendations of Ash et. al.. [56] by examining both error metrics — like Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) — and correlation metrics, including the coefficient of determination ($R^{2}$) and Pearson correlation ($\rho$). Error metrics assess the model’s accuracy in predicting absolute values, while correlation metrics evaluate its ability to capture trends within the data distribution. Ash et. al.. also advocate for parametric significance testing using repeated measures ANOVA (ANOVA-RM) followed by the Tukey Honestly Significant Difference (TukeyHSD) test. This approach accounts for the repeated-measures setup typical in our domain (i.e., evaluating different models on the same test set) and corrects for multiple comparisons across more than two models. The authors of Ash et. al. adapted the TukeyHSD implementation to use the standard error of the means derived from ANOVA-RM, rather than from a standard ANOVA. For pairwise model comparisons, we report statistical significance using paired t-tests. In the main manuscript, we primarily report MAE and $R^{2}$ as our core evaluation metrics; however, additional metrics such as RMSE, $\rho$, and Spearman correlation are available on our GitHub repository.
@@ 3.6 Models benchmarking
To assess the strengths and limitations of our transformer model, we compare its performance on downstream molecular property prediction tasks using embeddings extracted from our model, two state-of-the-art transformer models — MolBERT [20] and MolFormer [26] — and molecular features, i.e., physicochemical descriptors and Morgan fingerprints (default settings) computed using RDKit [52].
We used the original MolBERT implementation [20] and the HuggingFace implementation of MolFormer [57] to extract embeddings. A Random Forest (RF) model with default settings from scikit-learn [51] was trained to predict molecular properties based on each representation: our model’s embeddings, MolBERT embeddings, MolFormer embeddings, and molecular features.
Table 1 provides an overview of each model, including pre-training dataset size, training objectives, domain adaptation details, and the number of parameters. The baseline models serve as simple, transparent benchmarks for comparison against the more complex transformer models.
@@
RF + Morgan fingerprint - - - - -
Table 1: An overview of the benchmarked models explaining their set-up in terms of pre-training dataset size, objectives and number of trainable parameters. MLM: masked language modeling, MTR: multi-task regression of physicochemical properties, SMILES-Eq: predicting whether two SMILES strings correspond to the same molecule, and CL: contrastive learning of the different representations of the SMILES string.